# ATLAS Photometry of Interstellar Object 3I/ATLAS

### v1

@@ 3 Discussion and Conclusions
The evolution of the absolute $H$ brightness of 3I/ATLAS is visible in the right panel of Figure 3, although the two earliest epochs may not be fully reliable, as described above. The $H$ magnitudes at 60792 and 60828 are distinctly fainter than the later points, contrary to the ZTF magnitude reported by MPEC 2025-N51222https://minorplanetcenter.net/mpec/K25/K25N51.html, converted to $H_{V}$ by Seligman et al. (2025). Since neither uncertainty nor aperture were reported for the ZTF magnitude, this may not be a real discrepancy.
@@
  • The early light curve (MJD<60860<60860) appears quite red, $(o-c)\approx 0.7$, and fairly flat in $H$ magnitude prior to MJD 60845, which may indicate that material from the red surface of the nucleus (which evolved under a billion years of 50 km/s proton impact) was a significant contributor to the light. This is not to imply the nucleus has $H\approx 13$ but that, like Bennu, bits of surface may be ejected by CO2 pressure to form a coma, and that they are evaporating nearly as fast as they are being thrown off.
  • Between 60845<<MJD<<60860 the color remained red and the $H$ magnitude brightened significantly and then declined. This is the time of the discovery and the CFHT and VLT observations described in Seligman et al. (2025) of a solar-pointing “rubble fountain”, again possibly bits of red surface being ejected by CO2.
  • Since 60870 the color has been $(c-o)\approx 0.3$, close to the solar value of (co)=0.286(c-o)=0.286, perhaps indicating that the red particles from the crust are now being overwhelmed by white grains of ice.
  • The slope of $H(t)$ changed markedly around 60890 at heliocentric distance $\sim$3.3 au, from 0.035-0.035 mag/day to 0.014-0.014 mag/day, just about the time that 3I/ATLAS developed the anti-solar tail visible in Figure 2. Expressed in terms of $r$, this corresponds to a coma cross section that goes as $r^{-2}$ overall, but $r^{-3.9}$ prior to 60890 and $r^{-1.2}$ ever since. We speculate that this may be caused by enough tiny ice particles being generated to be pushed back in the anti-solar direction, and at the same time increasing the optical depth through the coma above unity so that a fraction of new mass being evolved is shadowed or not visible from Earth. (Or perhaps small ice grains are simply evaporating more quickly.)
This speculation only serves to highlight the features visible in the photometric data, and will be superseded by physical models of how 3I/ATLAS is really evolving.

### v3

@@ 3 Discussion and Conclusions
The evolution of the absolute $H$ brightness of 3I/ATLAS is visible in the right panel of Figure 3, although the two earliest epochs may not be fully reliable, as described above. The $H$ magnitudes at 60792 and 60828 are distinctly fainter than the later points, contrary to the ZTF magnitude reported by MPEC 2025-N51333https://minorplanetcenter.net/mpec/K25/K25N51.html, converted to $H_{V}$ by Seligman et al. (2025). Since neither uncertainty nor aperture were reported for the ZTF magnitude, this may not be a real discrepancy.
@@
  • The early light curve (MJD<60860<60860) appears quite red, $(o-c)\approx 0.7$, and fairly flat in $H$ magnitude for two months prior to MJD 60845.
  • Between 60845<<MJD<<60860 the color remained red and the $H$ magnitude brightened significantly and then declined. This is the time of the discovery and the CFHT and VLT observations described in Seligman et al. (2025) of a solar-pointing “rubble fountain”.
  • The continuous, 12 hour sequence of observation on 60861 shows a steady decline in brightness by +0.27 mag/day and RMS residual of 0.025 mag that is consistent with the uncertainties; no significant rotational modulation is visible.
  • Since 60870 the color has been $(c-o)\approx 0.3$, close to the solar value of (co)=0.286(c-o)=0.286.
  • The slope of $H(t)$ changed markedly around 60890 at heliocentric distance $\sim$3.3 au, from 0.035-0.035 mag/day to 0.012-0.012 mag/day, just about the time that 3I/ATLAS developed the anti-solar tail visible in Figure 2. Expressed in terms of $r$, this corresponds to a coma cross section that goes as $r^{-2}$ overall, but $r^{-3.9}$ prior to 60890 and $r^{-1.1}$ ever since.